# Probing scalar non-standard interaction of supernova neutrinos in next-generation neutrino experiments

### v1

@@ Abstract
A new neutrino-matter interaction can potentially affect neutrino propagation through matter. In this work, we explore the impact of a flavor-conserving scalar-mediated non-standard neutrino interaction in the supernova neutrino flux. We observe that the presence of scalar interaction involving muon and tau neutrinos (parameterized as $\eta_{\mu\mu}$ and $\eta_{\tau\tau}$, respectively) can invert the neutrino mass eigenstate in which three neutrino flavor states are produced inside the Supernova core, resulting in a significant modification of the electron neutrino flux from supernova reaching the Earth. In the context of the DUNE experiment, we estimate the number of supernova neutrino events in the presence of scalar non-standard neutrino interaction $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ and contrast with the case without scalar-mediated non-standard interactions. Our results indicate that such scalar interactions introduce a new degeneracy in the measurement of neutrino mass ordering from supernova neutrinos. We show how the $\bar{\nu}_{e}$ event distribution in Hyper-Kamiokande experiment can help resolve the degeneracy between a model with new scalar interactions for normal ordered neutrino masses and the standard model with inverted mass ordering.
@@ 1 Introduction
Inside the core of a core-collapse supernova, an enormous number of neutrinos ($\sim 10^{58}$) are generated through several processes in tens of seconds, carrying away 99% of the total emitted energy Janka:2017vlw ; Mirizzi:2015eza . These neutrinos experience highly dense and rapidly varying matter densities while escaping the supernova, which drastically influence the neutrino flavor evolution. Due to standard neutrino–matter interactions, active neutrinos, particularly electron neutrinos, undergo multiple adiabatic flavor transitions as they propagate from the dense core to the stellar envelope. These transitions are sensitive to the neutrino mass ordering Dasgupta:2008my that affects the outgoing electron neutrino flux. As a result, the neutrino signal from a future supernova explosion may play a crucial role in establishing the neutrino mass ordering. However, physics beyond the Standard Model (BSM) can potentially modify the picture Nunokawa:1997ct ; Esteban-Pretel:2007zkv ; Dighe:2017sur ; deGouvea:2019goq ; Chen:2022kal ; Tang:2020pkp ; Jana:2022tsa ; Jana:2023ufy ; dosSantos:2023skk ; Carenza:2023old ; Sen:2024fxa ; Jana:2024lfm . In this work, we investigate the impact of scalar-mediated non-standard interactions (SNSI) between neutrinos and background matter on supernova neutrino flavor evolution. These interactions modify the general Dirac mass term for neutrinos and can alter their propagation through matter Ge:2018uhz . SNSI effects have previously been explored in the context of solar and accelerator neutrino experiments Babu:2019iml ; Medhi:2021wxj ; Singha:2023set ; ESSnuSB:2023lbg ; Sarker:2023qzp ; Denton:2024upc ; Dutta:2024hqq . This work investigate the impact of flavor conserving SNSI on the supernova neutrino events in the upcoming neutrino experiments Deep Underground Neutrino Experiment (DUNE) and Hyper-Kamiokande (Hyper-K).
@@
DUNE DUNE:2020jqi and Hyper-K Hyper-Kamiokande:2018ofw are among the leading next-generation neutrino experiments. DUNE consists of a Liquid Argon Time Projection Chamber (LArTPC) of 40 kt located at the Sanford Underground Research Facility (SURF) in the USA. It will receive a muon neutrino beam from the accelerator at Fermilab, situated at a distance of approximately 1300 km from SURF. The primary goal of this experiment is to resolve the unknown issues in neutrino oscillation such as leptonic CP violation and neutrino mass ordering DUNE:2020jqi . In addition to its accelerator-based program, DUNE is capable of detecting neutrinos from core-collapse supernovae, thanks to its low energy threshold ($\sim 5$ MeV) DUNE:2020ypp . The large volume of the detector and excellent efficiency make it a promising observatory for capturing supernova neutrino signals. Hyper-K is the successor of the SuperKamiokande detector based in the Kamioka mine in Japan. It will eventually consist of two identical water Cherenkov detectors of 187 kt each, which includes a proposed second detector complex based in Korea Hyper-Kamiokande:2018ofw . The energy threshold to detect electron antineutrino events is as small as several MeV, making it ideal for supernova neutrino observations Abe:2011ts . We estimate the expected supernova neutrino event rate these two detectors would observe from future core-collapse supernovae and how the possible scalar-mediated interaction influences the event distributions.
@@
The paper is organized as follows: In section 2, we present the theoretical framework of scalar-mediated non-standard interactions and describe how these interactions modify the neutrino propagation Hamiltonian. Section 3 demonstrates the flavor evolution of the neutrino propagating from the supernova core to the surface and how the presence of scalar-mediated interaction modifies the propagation. In Section 4, we compute the expected event distributions of supernova neutrinos at DUNE and Hyper-K and discuss our results. Finally, we conclude in section 5.
@@ 2 Theoretical formalism of SNSI
where $y$ and $Y_{\alpha\beta}$ shows the coupling strength of charged fermions $f$ ($f=e,u,d$) and SM neutrinos $(\nu_{\alpha}=\nu_{e},\nu_{\mu},\nu_{\tau})$ with the scalar mediator $\phi$, respectively. The mediator mass $m_{\phi}$ and the couplings $Y_{\alpha\beta}$, $y_{f}$ are constrained by various observables like leptonic decay of mesons Pasquini:2015fjv , big bang nucleosynthesis (BBN) Huang:2017egl ; Babu:2019iml , coherent scattering Farzan:2018gtr , neutrino events from 1987A Heurtier:2016otg , fifth force experiments Babu:2019iml , and many others. This work focuses on scalar-mediated non-standard interactions of supernova neutrinos, where we consider the mediator mass to be less than 5 MeV Smirnov:2019cae ; Babu:2019iml .
@@
when non-relativistic fermions with number densities $n_{f}$ dominate effective mass correction, which we assume here. Note that, inside the supernova core, electrons in the background medium can be relativistic, for which the correction term is proportional to $n_{e}^{2/3}$ Babu:2019iml . However, since the scalar-mediated interactions are primarily governed by the nucleons inside the core, considering relativistic electrons does change our results qualitatively. The effective neutrino propagation Hamiltonian in the presence of SNSI is written as
@@
where $M_{\rm re}=M+\delta\tilde{M}$, $\rho$ is the local matter density converted from the fermion number density, and $\rho_{s}$ is a fixed scaling density, generally taken to be the average density of the medium. $\delta\tilde{M}$ denotes the matrix $\delta M$ computed at the scaling density $\rho_{s}$, such that $\delta M=\delta\tilde{M}\frac{\rho}{\rho_{s}}$. The matrix, $\delta\tilde{M}$, can be parameterized in terms of dimensionless quantity $\eta_{\alpha\beta}$ after scaling it with $\sqrt{\Delta m^{2}_{31}}\equiv\sqrt{m^{2}_{3}-m^{2}_{1}}$,
@@
where $\eta_{\alpha\beta}\propto\frac{\sum_{f}n_{f,s}y_{f}Y_{\alpha\beta}}{m^{2}_{\phi}\sqrt{\Delta m^{2}_{31}}}$, with $n_{f,s}$ denoting the average number density of the fermion $f$.
@@ 3 Neutrino flavor evolution in core-collapse supernova in the presence of flavor conserving SNSI
During the various stages of a core-collapse supernova, all three neutrino flavors ($\nu_{e}$, $\nu_{\mu}$, and $\nu_{\tau}$) are generated through different physical processes (see, e.g., ref. Janka:2017vlw ). During the neutronization-burst phase, which occurs within tens of milliseconds after the core bounce, electrons in the medium are captured by the free protons to produce a large number of electron neutrinos. As a result, the $\nu_{e}$ flux dominates over the other two SM neutrino flavors in this stage. The generated $\nu_{e}$ flux remains trapped in highly dense post-shock matter before they freely stream outwards once the surrounding matter becomes sufficiently transparent as the density drops Janka:2017vlw . Shortly after the neutronization burst, the escape of electron neutrinos reduces the electron lepton number in the medium. Consequently, the concentration of positrons rises, enhances electron-positron pair production, and leads to the production of the other two flavors of neutrinos and antineutrinos ($\nu_{x},\bar{\nu}_{x}=\nu_{\mu},\nu_{\tau},\bar{\nu}_{\mu},\bar{\nu}_{\tau}$). Additionally, positrons capture neutrons to produce electron antineutrinos.
The time and energy-dependent distribution of the unoscillated neutrino flux emitted during a supernova explosion can be expressed as
@@
The neutrino flux $\phi(E,t)$ generated within the supernova core for each flavor is significantly modified by the neutrino flavor mixing effect. In the dense inner regions of the exploding star, where matter densities can reach up to $\sim 10^{14}$ g/cm3, neutrino-matter interactions—particularly the CC interactions involving electron neutrinos—play a dominant role in the neutrino mixing, thus affecting the neutrino flavor evolution. In addition, possible scalar-mediated non-standard interactions can further influence the neutrino flavor evolution. As neutrinos propagate through the varying matter density of the stellar environment, the effective mixing between the neutrino flavor states varies with local matter density and neutrino energy. To have a better understanding of the effective neutrino mixing inside the star during a supernova explosion, we diagonalize the total neutrino propagation Hamiltonian, as defined in Eq. (6), $\mathcal{H}_{\rm diag}={U^{\rm eff}}^{\dagger}\cdot\mathcal{H}\cdot U^{\rm eff}$. The effective mixing matrix $U^{\rm eff}$ determines the mixing between the neutrino flavor states and the effective mass states, $\nu_{i,m}$ ($i=1,2,3$).
@@
We consider a radial matter density profile that decreases from $\sim 10^{14}$ g/cm3 at the stellar core to zero at the envelope Tang:2020pkp ; Jana:2024lfm , assuming an average matter density $\rho_{s}=10^{6}$ g/cm3. The matter potential is computed using the electron number density profile described in Refs. Tang:2020pkp ; Jana:2024lfm . In the standard model case, where outgoing neutrinos interact with the highly dense matter ($\rho>10^{5}$ g/cm3) inside the core via CC and NC interactions, the effective mixing matrix elements for electron neutrinos take the values $|U^{\rm eff}_{e3}|\simeq 1$, $|U^{\rm eff}_{e2}|\simeq|U^{\rm eff}_{e1}|\simeq 0$ for normal mass ordering (NMO) of neutrinos ($m_{1}<m_{2}<m_{3})$. In other words, electron neutrino flux produced inside the supernova primarily aligns with the $\nu_{3,m}$ states. It is assumed that the propagation of the neutrino mass states through the varying matter density in the supernova maintains adiabaticity and is incoherent, such that $(\nu_{i,m})$ states reach the surface of the supernova as the $\nu_{i}$ state. This means that the flux of $\nu_{e}$, $\Phi(\nu_{e})\simeq\Phi(\nu_{3,m})$ reaches the surface as $\Phi(\nu_{3})$. The resulting expression for the $\Phi(\nu_{e})$ flux at the supernova surface, assuming NMO in the standard model, is given by
$\Phi(\nu_{e})=|U_{e3}|^{2}\Phi_{c}(\nu_{e})+(1-|U_{e3}|^{2})\,\Phi_{c}(\nu_{x}),$ (11)
@@
where $\Phi_{c}(\nu_{\alpha})$ denotes the $\nu_{\alpha}$ flux at the supernova core estimated using eq. (9). The last term in eq. (11) accounts for the contribution from the other two flavors of the neutrino, assuming $\Phi_{c}(\nu_{\mu})=\Phi_{c}(\nu_{\tau})=\Phi_{c}(\nu_{x})$. For the inverted mass ordering (IMO) of neutrino ($m_{3}<m_{1}<m_{2}$), $\nu_{e}$ aligns with the $\nu_{2,m}$ states, and the fraction of electron neutrinos at the supernova surface is
$\Phi(\nu_{e})=|U_{e2}|^{2}\Phi_{c}(\nu_{e})+(1-|U_{e2}|^{2})\Phi_{c}(\nu_{x})\,.$ (12)
@@
Figure 1: Effective mixing elements $|U_{e1}|$, $|U_{e2}|$, and $|U_{e3}|$ as a functions of matter density in the presence of scalar mediated NSI parameters, $\eta_{ee}$ (left column), $\eta_{\mu\mu}$ (middle column), and $\eta_{\tau\tau}$ (right colum). The top and bottom rows correspond to normal mass ordering and inverted mass ordering of neutrinos, respectively. We consider neutrino energy $E=10$ MeV in the plot. The value of oscillation parameters used to compute the mixing matrix elements are taken from NuFit-6 Esteban:2024eli .

### v2

@@ Abstract
A new neutrino-matter interaction can potentially affect neutrino propagation through matter. In this work, we explore the impact of a flavor-conserving scalar-mediated non-standard neutrino interaction in the supernova neutrino flux. We observe that the presence of a scalar interaction involving muon and tau neutrinos (parameterized as $\eta_{\mu\mu}$ and $\eta_{\tau\tau}$, respectively) can invert the neutrino mass eigenstate in which three neutrino flavor states are produced inside the supernova core, resulting in a significant modification of the electron neutrino flux from the supernova reaching the Earth. In the context of the DUNE experiment, we estimate the number of supernova neutrino events in the presence of scalar non-standard neutrino interaction $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ and contrast with the case without scalar-mediated non-standard interactions. Our results indicate that such scalar interactions introduce a new degeneracy in the measurement of neutrino mass ordering from supernova neutrinos. We show how the $\bar{\nu}_{e}$ event distribution in Hyper-Kamiokande experiment may help resolve the degeneracy between a model with new scalar interactions for normal ordered neutrino masses and the standard model with inverted mass ordering for a galactic supernova.
@@ 1 Introduction
Inside the core of a core-collapse supernova, an enormous number of neutrinos ($\sim 10^{58}$) are generated through several processes in tens of seconds, carrying away 99% of the total emitted energy Janka:2017vlw ; Mirizzi:2015eza . These neutrinos experience highly dense and rapidly varying matter densities while escaping the supernova, which drastically influence the neutrino flavor evolution. Due to standard neutrino–matter interactions, active neutrinos, particularly electron neutrinos, undergo multiple adiabatic flavor transitions as they propagate from the dense core to the stellar envelope Dighe:1999bi . These transitions are sensitive to the neutrino mass ordering Dasgupta:2008my that affects the outgoing electron neutrino flux. As a result, the neutrino signal from a future supernova explosion may play a crucial role in establishing the neutrino mass ordering. However, physics beyond the Standard Model (BSM) can potentially modify the picture Nunokawa:1997ct ; Esteban-Pretel:2007zkv ; Dighe:2017sur ; deGouvea:2019goq ; Chen:2022kal ; Tang:2020pkp ; Jana:2022tsa ; Jana:2023ufy ; dosSantos:2023skk ; Carenza:2023old ; Sen:2024fxa ; Jana:2024lfm . In this work, we investigate the impact of scalar-mediated non-standard interactions (SNSI) between neutrinos and background matter on supernova neutrino flavor evolution. These interactions modify the general Dirac mass term for neutrinos and can alter their propagation through matter Ge:2018uhz . SNSI effects have previously been explored in the context of solar and accelerator neutrino experiments Babu:2019iml ; Medhi:2021wxj ; Singha:2023set ; ESSnuSB:2023lbg ; Sarker:2023qzp ; Gupta:2023wct ; Denton:2024upc ; Dutta:2024hqq . This work investigate the impact of flavor conserving SNSI on the supernova neutrino events in the upcoming neutrino experiments Deep Underground Neutrino Experiment (DUNE) and Hyper-Kamiokande (Hyper-K).
@@
DUNE DUNE:2020jqi ; DUNE:2020ypp and Hyper-K Hyper-Kamiokande:2018ofw are among the leading next-generation neutrino experiments. DUNE consists of a Liquid Argon Time Projection Chamber (LArTPC) of 40 kt located at the Sanford Underground Research Facility (SURF) in the USA. It will receive a muon neutrino beam from the accelerator at Fermilab, situated at a distance of approximately 1300 km from SURF. The primary goal of this experiment is to resolve the unknown issues in neutrino oscillation such as leptonic CP violation and neutrino mass ordering DUNE:2020jqi . In addition to its accelerator-based program, DUNE is capable of detecting neutrinos from core-collapse supernovae, thanks to its low energy threshold ($\sim 5$ MeV) DUNE:2020ypp . The large volume of the detector and excellent efficiency make it a promising observatory for capturing supernova neutrino signals. Hyper-K is the successor of the Super-Kamiokande detector based in the Kamioka mine in Japan. It will eventually consist of two identical water Cherenkov detectors of 187 kt each, which includes a proposed second detector complex based in Korea Hyper-Kamiokande:2018ofw . The energy threshold to detect electron antineutrino events is as small as several MeV, making it ideal for supernova neutrino observations Abe:2011ts . We estimate the expected supernova neutrino event rate these two detectors would observe from future core-collapse supernovae and how the possible scalar-mediated interaction influences the event distributions.
@@
The paper is organized as follows: In section 2, we present the theoretical framework of scalar-mediated non-standard interactions and describe how these interactions modify the neutrino propagation Hamiltonian. Section 3 demonstrates the flavor evolution of the neutrino propagating from the supernova core to the surface and how the presence of scalar-mediated interaction modifies the propagation. In Section 4, we compute the expected event distributions of supernova neutrinos at DUNE and Hyper-K and discuss our results. Finally, we conclude in section 5. In appendix A, we show the adiabaticity condition for neutrino propagation in the presence of scalar NSI.
@@ 2 Theoretical formalism of SNSI
where $y$ and $Y_{\alpha\beta}$ show the coupling strength of charged fermions $f$ ($f=e,u,d$) and SM neutrinos $(\nu_{\alpha}=\nu_{e},\nu_{\mu},\nu_{\tau})$ with the scalar mediator $\phi$, respectively. The mediator mass $m_{\phi}$ and the couplings $Y_{\alpha\beta}$, $y_{f}$ are constrained by various observables like the leptonic decay of mesons Pasquini:2015fjv , big bang nucleosynthesis (BBN) Huang:2017egl ; Babu:2019iml , coherent scattering Farzan:2018gtr , neutrino events from 1987A Heurtier:2016otg ; Fiorillo:2022cdq , fifth force experiments Babu:2019iml , and many others. The expected constraints on $Y_{\alpha\beta}$ and $m_{\phi}$ from future supernova events have been discussed in refs. Akita:2022etk ; Lazar:2024ovc ; Telalovic:2024cot ; Akita:2023iwq .
@@
when non-relativistic fermions with number densities $n_{f}$ dominate the effective mass correction, which we assume here. Note that, inside the supernova core, electrons in the background medium can be relativistic, for which the correction term is proportional to $n_{e}^{2/3}$ Babu:2019iml . However, since the scalar-mediated interactions are primarily governed by the nucleons inside the core, considering relativistic electrons does not change our results qualitatively. Based on ref. Babu:2019iml , we estimate that the scalar NSI contribution from relativistic electrons in the supernova core is suppressed by almost three orders of magnitude compared to that from non-relativistic nucleons. The effective neutrino propagation Hamiltonian in the presence of SNSI is written as
@@
where $M_{\rm re}=M+\delta\tilde{M}$, $\rho$ is the local matter density converted from the fermion number density, and $\rho_{s}$ is a fixed scaling density, generally taken to be the average density of the medium. We use $\rho_{s}=10^{5}$ g/$c$. $\delta\tilde{M}$ denotes the matrix $\delta M$ computed at the scaling density $\rho_{s}$, such that $\delta M=\delta\tilde{M}\frac{\rho}{\rho_{s}}$. The matrix, $\delta\tilde{M}$, can be parameterized in terms of dimensionless quantities $\eta_{\alpha\beta}$ after scaling it with $\sqrt{|\Delta m^{2}_{31}|}\equiv\sqrt{|m^{2}_{3}-m^{2}_{1}|}$,
@@
where $\eta_{\alpha\beta}\propto\frac{\sum_{f}n_{f,s}y_{f}Y_{\alpha\beta}}{m^{2}_{\phi}\sqrt{|\Delta m^{2}_{31}|}}$, with $n_{f,s}$ denoting the average number density of the fermion $f$. This work focuses on scalar-mediated non-standard interactions of supernova neutrinos. We consider the SNSI correction term $\delta M$ to be less than 5 MeV Smirnov:2019cae ; Babu:2019iml . We also assume neutrino coupling $y_{\nu}$ to be small $y_{\alpha\beta}\lesssim 10^{-10}$, as a large coupling may lead to a significant contribution from neutrino self-interactions Fiorillo:2023cas ; Fiorillo:2023ytr .
@@ 3 Neutrino flavor evolution in core-collapse supernova in the presence of flavor conserving SNSI
During the various stages of a core-collapse supernova, all three neutrino flavors ($\nu_{e}$, $\nu_{\mu}$, and $\nu_{\tau}$) are generated through different physical processes (see, e.g., ref. Janka:2017vlw ). During the neutronization-burst phase, which occurs inside the core of the star during tens of milliseconds after the core bounce, electrons in the medium are captured by the free protons $(e^{-}+p\rightarrow n+\nu_{e})$ to produce a large number of electron neutrinos. As a result, the $\nu_{e}$ flux dominates over the other two SM neutrino flavors in this stage. The generated $\nu_{e}$ flux remains trapped in highly dense post-shock matter before they freely stream outwards once the surrounding matter becomes sufficiently transparent as the density drops Janka:2017vlw . Shortly after the neutronization burst, the loss of electron–lepton number due to the escape of the $\nu_{e}$ flux leads to the thermal production of positrons. This accelerates the production of $\bar{\nu}_{e}$ and the two other neutrino and antineutrino flavors ($\nu_{x},\bar{\nu}_{x}=\nu_{\mu},\nu_{\tau},\bar{\nu}_{\mu},\bar{\nu}_{\tau}$) via pair-annihilation ($e^{+}+e^{-}\rightarrow\nu+\bar{\nu}$) and nucleon-nucleon bremsstrahlung ($N+N^{\prime}\rightarrow N+N^{\prime}+\nu+\bar{\nu}$). Additionally, positron capture on neutrons ($e^{+}+n\rightarrow p+\bar{\nu}_{e}$) contributes to the production of $\bar{\nu}_{e}$ at the end of the neutronization phase and in the post-neutronization phase. In the post-neutronization phase, since there is a high density of neutrinos and comparatively low matter density due to the expansion of the shock wave, collective neutrino oscillations from neutrino self-interactions play an important role in the neutrino flavor evolution Duan:2010bg ; Chakraborty:2016yeg . However, in this work, we place a particular emphasis on the impact of neutrino-matter interactions in the flavor evolution, most important in the neutronization-burst phase when the effects from collective neutrino oscillations are relatively suppressed as compared to matter-induced flavor evolution in the high matter densities inside the core. A full treatment that also considers the impact of collective oscillations in neutrino propagation is beyond scope of this work. We note that the $\bar{\nu}_{e}$ flux in the neutronization phase is very small compared to the $\nu_{e}$ flux. The $\bar{\nu}_{e}$ flux may be large enough for detection by large volume experiments such as Hyper-K deGouvea:2019goq , but only if supernova neutrino luminosity is large enough, and the supernova is relatively close, as discussed below.
The time- and energy-dependent distribution of the unoscillated neutrino flux emitted during a supernova explosion can be expressed as
@@
The neutrino flux $\Phi_{\nu}(E,t)$ generated within the supernova core for each flavor is significantly modified by the neutrino flavor mixing effect. In the dense inner regions of the exploding star, where matter densities can reach up to $\sim 10^{14}$ g/cm3, neutrino-matter interactions—particularly the CC interactions involving electron neutrinos—play a dominant role in the neutrino mixing, thus affecting the neutrino flavor evolution. In addition, possible scalar-mediated non-standard interactions can further influence the neutrino flavor evolution. As neutrinos propagate through the varying matter density of the stellar environment, the effective mixing between the neutrino flavor states varies with local matter density and neutrino energy. To have a better understanding of the effective neutrino mixing inside the star during a supernova explosion, we diagonalize the total neutrino propagation Hamiltonian, as defined in eq. (6), $\mathcal{H}_{\rm diag}={U^{\rm eff}}^{\dagger}\cdot\mathcal{H}\cdot U^{\rm eff}$. The effective mixing matrix $U^{\rm eff}$ determines the mixing between the neutrino flavor states and the effective mass states, $\nu_{i,m}$ ($i=1,2,3$).
@@
We consider a radial matter density profile that decreases from $\sim 10^{14}$ g/cm3 at the stellar core to zero at the envelope Tang:2020pkp ; Jana:2024lfm . The matter potential is computed using the electron number density profile described in Refs. Tang:2020pkp ; Jana:2024lfm . In the standard model case, where outgoing neutrinos interact with the highly dense matter ($\rho>10^{5}$ g/cm3) inside the core via CC and NC interactions, the effective mixing matrix elements for electron neutrinos take the values $|U^{\rm eff}_{e3}|\simeq 1$, $|U^{\rm eff}_{e2}|\simeq|U^{\rm eff}_{e1}|\simeq 0$ for normal mass ordering (NMO) of neutrinos ($m_{1}<m_{2}<m_{3})$ Dighe:1999bi . In other words, electron neutrino flux produced inside the supernova primarily aligns with the $\nu_{3,m}$ state. It is verified that the propagation of the neutrino mass states through the varying matter density in the supernova maintains adiabaticity and is incoherent, such that $(\nu_{i,m})$ state reaches the surface of the supernova as the $\nu_{i}$ state. This means that the flux of $\nu_{e}$, $\Phi(\nu_{e})\simeq\Phi(\nu_{3,m})$ reaches the surface as $\Phi(\nu_{3})$. The resulting expression for the $\Phi(\nu_{e})$ flux at the supernova surface, assuming NMO in the standard model (SI), is given by
$\Phi(\nu_{e})=|U_{e3}|^{2}\Phi_{c}(\nu_{e})+(1-|U_{e3}|^{2})\,\Phi_{c}(\nu_{x})\,\,\,\,\,\,...\,\,{\rm for\ NMO(SI)},$ (11)
@@
where $\Phi_{c}(\nu_{\alpha})$ denotes the $\nu_{\alpha}$ flux at the supernova core estimated using eq. (9). The last term in eq. (11) accounts for the contribution from the other two flavors of the neutrino, assuming $\Phi_{c}(\nu_{\mu})=\Phi_{c}(\nu_{\tau})=\Phi_{c}(\nu_{x})$. For the inverted mass ordering (IMO) of neutrino ($m_{3}<m_{1}<m_{2}$), $\nu_{e}$ aligns with the $\nu_{2,m}$ state, and the fraction of electron neutrinos at the supernova surface is
$\Phi(\nu_{e})=|U_{e2}|^{2}\Phi_{c}(\nu_{e})+(1-|U_{e2}|^{2})\Phi_{c}(\nu_{x})\,\,\,\,\,\,...\,\,{\rm for\ IMO(SI)}\,.$ (12)
@@
Figure 1: For normal mass ordering, the effective mixing elements $|U_{e1}^{\rm eff}|$, $|U_{e2}^{\rm eff}|$, and $|U_{e3}^{\rm eff}|$ as a functions of matter density in the presence of scalar mediated NSI parameters, $\eta_{ee}$ (left column), $\eta_{\mu\mu}$ (middle column), and $\eta_{\tau\tau}$ (right column). The top and bottom rows correspond to the strength of the SNSI parameters $\eta_{\alpha\alpha}=0.1$ and $\eta_{\alpha\alpha}=0.01$, respectively. We consider neutrino energy $E=10$ MeV in the plot. The value of oscillation parameters used to compute the mixing matrix elements are taken from NuFit-6 Esteban:2024eli .
Figure 2: Same as fig. 1, but with inverted mass ordering of neutrino masses.